OpenStudy (anonymous):

A circular sector with radius r and angle Q has area A. Find r and Q so that the perimeter is smallest for a given area A. (Note: A=1/2r^2Q, and the length of the arc s=rQ, when Q is measured in radians)?

OpenStudy (dumbcow):

I believe its r = sqrt(A) and Q = 2 is this for a calculus class?

OpenStudy (anonymous):

Yes. how did you get it?

OpenStudy (dumbcow):

ok just wanted to make sure because it involves taking a derivative to minimize the perimeter. \[A = \frac{r^{2}Q}{2} , P = 2r + rQ\] Now using the Area equation, solve for Q \[Q = \frac{2A}{r^{2}}\] Substitute into Perimeter equation \[P = 2r + r(\frac{2A}{r^{2}}) = 2r + \frac{2A}{r}\] Differentiate \[dP/dr = 2 - \frac{2A}{r^{2}}\] Set equal to zero and solve for r \[2 - \frac{2A}{r^{2}} = 0\] \[\rightarrow 2r^{2} = 2A\] \[\rightarrow r = \sqrt{A}\] Now substitute this value back in for r to solve for Q. \[Q = \frac{2A}{r^{2}} = \frac{2A}{\sqrt{A}^{2}} = \frac{2A}{A} = 2\]

OpenStudy (anonymous):

Thank you so much. But can you tell me how did you compute P=2r+rQ? Is this a formula?

OpenStudy (dumbcow):

no problem its the sum of the arc length and the 2 radius. a sector is like an isosceles triangle except the base is curved but its length is the arc length

OpenStudy (anonymous):

Okay, thanks. Can you please go through more word problems. I have a exam tomorrow so I really need to figure these out!